Suppose a viscous oil, whose flow is in the laminar regime is to be pumped through a 10 cm diameter horizontal pipe over a distance of 15km at a rate of 10^{-3} m^3/s. Viscosity of the oil is 0.03 poise. What is the required pressure drop to maintain such a flow?
1
Expert's answer
2020-06-17T09:42:42-0400

The number of units of fluid flowing per unit time through the cross section of the pipe:

"Q= \frac{\pi R^4} {8 \eta l} \Delta P" , where:

"R" - radius of pipe, "R = 0.05m" ;

"\eta" -  viscosity of the oil, "\eta=0.03 poise" ;

"l" -pipe length, "l=15000m;"

"\Delta P" - the required pressure.

"\Delta P= \frac{Q}{\pi R^4}8\eta l = \frac{10^{-3}}{3.14\cdot 0.05^4}8\cdot0.03\cdot15000 = 183439.5 Pa"